Solve the given differential equation by using an appropriate substitution. The DE is a Bernoulli equation.




dy/dx = y(xy^3-1)

1
Expert's answer
2021-11-09T08:48:15-0500

Solution

Let’s rewrite equation in the form

dy/dx + y = xy^4

New function u(x) = y-3   

du/dx = -3 y-4 (dy/dx)   => dy/dx = - (du/dx) y4 /3

Substitution in equation gives new equation for u:

du/dx – 3u = –3x

It is linear equation . So  u(x) = C e3x + Ax + B, where A, B, C – are arbitrary constants,  C e3x – is general solution of homogeneous  equation.

Substitution into equation on u gives:

A – 3B – 3Ax = -3x => A = 1, B = 1/3.

Therefore u(x) = C e3x + x + 1/3 and

"y(x)=\sqrt[3]{\frac{1}{u(x)}}=\frac{1}{\sqrt[3]{Ce^{3x}+x+1/3}}"